A student prepared for the exam only one ticket №10 from 25 given ones. In what
turn it will be better for him to answer: the first, the second or the third?
1
Expert's answer
2020-11-05T17:44:11-0500

A - student gets ticket №10

If in the first turn:

p(A) = 1/25 = 0.04

If in the second turn:

p(A) = p(probability of getting ticket №10 in case the first student doesn't get ticket №10)*p(the first student doesn't get ticket №10)

p(A) = 1/24*24/25 = 1/25 = 0.04

If in the third turn:

p(A) = p(probability of getting ticket №10 in case the first and the second students don't get ticket №10)*p(the first student doesn't get ticket №10)*p(the second student doesn't get ticket №10)

p(A) = 1/23*24/25*23/24 = 1/25 = 0.04

Answer: it does not matter